OpenStudy (anonymous):

14y + 20 = 4y A. –2 B. 2 C. no solution D. all real numbers

OpenStudy (agent47):

-2

OpenStudy (agent47):

14y+20=4y 14y+20-4y=4y-4y 14y-4y+20=0 10y+20-20=0-20 10y=-20 y=-20/10=-2

OpenStudy (anonymous):

a. -2